OpenStudy (anonymous):

Find the exact values of sin(2θ), cos(2θ), and tan(2θ) for the given value of θ. sec(θ)=-3/2 ; 90<θ<180 sin(2θ) = cos(2θ) = tan(2θ) = how the heck am i gonna solve this..?

OpenStudy (camper4834):

find theta with sec(theta) = -3/2

OpenStudy (anonymous):

huh.. i didnt get it..?

OpenStudy (camper4834):

do you know that sec is 1/cos?

OpenStudy (anonymous):

reciprocal identity..

OpenStudy (anonymous):

so it will be -2/3..?

OpenStudy (camper4834):

yes cos = that

OpenStudy (anonymous):

so i will use double angle identity..?

OpenStudy (camper4834):

what? no

OpenStudy (anonymous):

i dont really get it.. am i just gonna plug it in or what..?

OpenStudy (camper4834):

oh i see you meant for the sin(2theta) yes you will use double angle

OpenStudy (camper4834):

also are you sure the question doesn't say sec(theta) = -sqrt(3)/2?

OpenStudy (anonymous):

sigh* i thought i was wrong.. hahah.. ok.. double angle just plug it or need one more to solve the exact value..?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so the cos(2theta) = -1/9..?

OpenStudy (camper4834):

yes

OpenStudy (anonymous):

how am i suppose to solve tan(2theta) where it is equal to 2tan(theta)/1-tan(theta)..?

OpenStudy (anonymous):

or because tan is equal to sin/cos

OpenStudy (camper4834):

lets start with sin

OpenStudy (camper4834):

sin(arccos(x) = sqrt(1-x^2)

OpenStudy (anonymous):

ow yeah.. the sin.. hahah.. forgot about that.. how am i gonna solve for sin it is equal to sin(2theta)=2sin(theta)cos(theta).. sin is missing and sin is needed to solve sin..? mind blown..

OpenStudy (anonymous):

whoa whoa whoa what is arccos..?

OpenStudy (camper4834):

\[\cos^{-1} \]

OpenStudy (anonymous):

ahhh.. ok hahha.. thats arcos.. raised to -1

OpenStudy (camper4834):

so knowing that \[\sin (\cos^{-1} (x) = \sqrt(1-x^2)\]

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

to solve x i need to get x y r..?

OpenStudy (camper4834):

\[\sin (2 \cos^{-1} (-2/3))=\sqrt(1-(-2/3)^2)*\cos(\cos^{-1} (-2/3))\]

OpenStudy (anonymous):

now it gets complicated..

OpenStudy (camper4834):

no no btw slow computer so sorry but theta = cos^-1(-2/3) right?

OpenStudy (anonymous):

yep

OpenStudy (camper4834):

okay so \[\sin (2\theta)=2\sin (\theta)\cos(\theta)\]

OpenStudy (anonymous):

so it is ewual to -2 square root of 5/9

OpenStudy (camper4834):

substitute cos^-1(-2/3) with theta

OpenStudy (anonymous):

i mean \[-\frac{ 2\sqrt{5} }{ 9 }\]

OpenStudy (camper4834):

almost i forgot the 2 in 2sinxcosx!

OpenStudy (anonymous):

so u mean theta is equal to cos^-1(-2/3)..?

OpenStudy (camper4834):

so your answer times two! and you are great!

OpenStudy (camper4834):

yes

OpenStudy (camper4834):

how did you solve the cos(2x) then?

OpenStudy (anonymous):

this one..? cos(2theta) = -1/9..

OpenStudy (camper4834):

yes how did you solve it without cos^-1

OpenStudy (anonymous):

coz cos(2theta) is also equal to 2cos(theta)^2-1

OpenStudy (camper4834):

XD oh you don't need cos^-1 for that....

OpenStudy (camper4834):

okay so you have sin and cos now for tangent!

OpenStudy (camper4834):

solve sin/cos!

OpenStudy (anonymous):

weyt u were way to fast.. were are we now..? haha

OpenStudy (camper4834):

\[\frac{ \sqrt(1-(-\frac{ 2 }{ 3 }) }{ -2/3 }\]

OpenStudy (camper4834):

oh im sorry, remember we did sin it was that crazy number \[\frac{ -4\sqrt(5) }{ 9 }\]

OpenStudy (anonymous):

\[-\frac{ \sqrt{15} }{ 2 }\] so is this it..?

OpenStudy (camper4834):

not 15 just 5

OpenStudy (anonymous):

my calculator says its 15..

OpenStudy (camper4834):

....i keep screwing up the formula

OpenStudy (anonymous):

what..? hahah.. and by the way its -2 not -4.. haha

OpenStudy (anonymous):

\[-\frac{ 2\sqrt{5} }{ 9 }\] thats it not -4

OpenStudy (camper4834):

again i screwed up the formula that time too i told you there was an extra TIMES 2 that i forgot

OpenStudy (anonymous):

hahaha.. now im confused.. hahaha

OpenStudy (anonymous):

so u mean i need to multiply it to 2 or not..?

OpenStudy (camper4834):

take your \[\frac{ -2\sqrt5 }{ 9 }\] and multiply it by 2

OpenStudy (camper4834):

you only solved sin(2x) = sin(x)cos(x) but it should be 2 sin(x)cos(x)

OpenStudy (camper4834):

nevermind just trust me

OpenStudy (anonymous):

so u said that u forgot to say that i need it to multiply to 2..? hHAHAHAH.. i know i trust u.. hahaha ok where are we.. on the -4 which is it is the sin..?

OpenStudy (camper4834):

if you want reinput these formulas sin(2x)=\[2\sqrt{1-\frac{ -2 }{ 3 }}*-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

ok so that is sin..?

OpenStudy (anonymous):

and u are right it is -4

OpenStudy (camper4834):

here is where we are \[\sin(2x)= \frac{ -4\sqrt5 }{ 9 }\] \[\cos(2x) = -\frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

how about the \[-\frac{ \sqrt{15} }{ 9 }\] is it 15 or 5..?

OpenStudy (camper4834):

okay let me put in the correct one this time

OpenStudy (camper4834):

\[\frac{ \sqrt{1-(\frac{ -2 }{ 3 })^2} }{ 2/3 }\]

OpenStudy (camper4834):

equals \[-\frac{ \sqrt5 }{ 2 }\]

OpenStudy (camper4834):

this is TAN(X) not TAN(2X) so all you need to do is what?

OpenStudy (camper4834):

\[\frac{ 2\tan(x) }{ 1-\tan(x)^2 }\]

OpenStudy (anonymous):

x is -sqrt 5/2..?

OpenStudy (anonymous):

or solve tan then plug it in that equation..?

OpenStudy (camper4834):

yes yes but guess what!!

OpenStudy (anonymous):

what..?

OpenStudy (camper4834):

we just now did that!

OpenStudy (anonymous):

whoa..? really how..?

OpenStudy (camper4834):

tan is \[-\sqrt5/2\]

OpenStudy (anonymous):

hwahaha.. that was unexpected.. hahaha..

OpenStudy (camper4834):

that crazy formula was sin/cos

OpenStudy (camper4834):

i just reformed it

OpenStudy (anonymous):

hey can i ask a favor..?

OpenStudy (camper4834):

sure

OpenStudy (anonymous):

from what country are u..?

OpenStudy (camper4834):

merica

OpenStudy (camper4834):

why

OpenStudy (anonymous):

im from philippines.. can u be online tomorrow..? what time is it in your america..?

OpenStudy (camper4834):

2:25AM

OpenStudy (camper4834):

crazy that im still up huh?

OpenStudy (anonymous):

oh um.. can u be online earlier like 12 am..? gamer peeps..?

OpenStudy (camper4834):

12 am my time?

OpenStudy (anonymous):

yep is that ok..? in our time its like 4 pm..

OpenStudy (camper4834):

yea sure XD

OpenStudy (anonymous):

yey.. coz i need to log out now.. see u tomorrow i need help about that again tomorrow.. by the way.. thanks.. im expecting u tom.. thanks again.. :)

OpenStudy (anonymous):

bye

OpenStudy (camper4834):

bye